# Reducing Biases in Record Matching Through Scores Calibration

### v2

@@ Abstract
Record matching is the task of identifying records that refer to the same real-world entity across datasets. While most existing models optimize for accuracy, fairness has become an important concern due to the potential for unequal outcomes across demographic groups. Prior work typically focuses on binary outcomes evaluated at fixed decision thresholds. However, such evaluations can miss biases in matching scores–biases that persist across thresholds and affect downstream tasks.
We propose a threshold-independent framework for measuring and reducing score bias, defined as disparities in the distribution of matching scores across groups. We show that several state-of-the-art matching methods exhibit substantial score bias, even when appearing fair under standard threshold-based metrics. To address this, we introduce two post-processing score calibration algorithms. The first, Calib, aligns group-wise score distributions using the Wasserstein barycenter, targeting demographic parity. The second, C-Calib, conditions on predicted labels to further reduce label-dependent biases, such as equal opportunity.
Both methods are model-agnostic and require no access to model training data. Calib also offers theoretical guarantees, ensuring reduced bias with minimal deviation from original scores. Experiments across real-world datasets and matching models confirm that Calib and C-Calib substantially reduce score bias while minimally impacting model accuracy.
Index Terms:
Record Matching, Fairness, Wasserstein Barycenter, Calibration, Threshold Independent.
I Introduction
@@
From a technical perspective, matching scores support the development of more flexible and advanced methods. For example, active learning approaches can use scores to identify uncertain cases—those with values near the decision threshold—and prioritize them for manual labeling to improve the model [1, 2]. This idea also extends to human-in-the-loop systems, where scores help determine which record pairs should be reviewed by a human. Instead of treating all pairs equally, the system can focus on those that are harder to classify automatically [3, 4]. Another technical reason for using scores instead of binary outcome is that many state-of-the-art matching techniques, especially deep learning models, naturally produce scores through their architecture—for example, by applying a sigmoid activation to the output of a final linear layer, resulting in a value between 0 and 1 that reflects the predicted match likelihood. Because of these advantages, matching scores are more widely used than hard binary decisions in both research and practical systems.
@@
Example 1.
Figure 1b presents another view of score bias by showing the ROC curves for the two groups. These curves measure the trade-off between TPR and false positive rate (FPR), and their area under the curve (AUC) is often used to assess score quality independent of any specific threshold. Although the ROC curves differ, the AUC values are nearly the same, as the differences cancel out. This shows that aggregate measures like AUC—despite being threshold-independent—can still overlook important biases in score distributions. $\blacksquare$
@@
Fairness in matching scores is also related to fairness in regression [14, 15, 16] and ranking [17, 18, 19, 20, 21, 22]. Like regression, matching scores are continuous, and we adopt a similar strategy of comparing score distributions across groups. However, unlike typical regression outputs, which often span wide or unbounded ranges, matching scores are usually concentrated near 0 and 1. We exploit this property in our bias reduction algorithms.
Fair ranking is also relevant, as both settings involve scoring and ranking inputs. However, fair ranking methods (e.g., [20, 21, 22]) focus on group-level balance among top-ranked items (e.g., top-$k$). In contrast, fairness in matching scores aims to ensure equitable outcomes across all score thresholds.
@@
To address fairness in matching scores, our first contribution is a new bias measure that extends standard notions such as demographic parity (DP) and equalized odds (EOD) from binary outcomes to score distributions across all thresholds. It captures cumulative performance differences between groups over the full threshold range. For example, in Figure 1a, the DP score bias is the sum of TPR differences across thresholds, shown by the gray area between the two plots. Using this measure, we evaluate popular benchmarks and state-of-the-art matching methods, revealing substantial score biases.
To reduce these biases, we propose a post-processing algorithm that calibrates scores without altering matching models, making it broadly applicable. The algorithm treats the scores for each demographic group as probability distributions and aligns them by mapping both to a shared target distribution. We use optimal transport (OT) to compute the Wasserstein barycenter [15, 23], which serves as a balanced average that minimizes deviation from the original scores while improving fairness. This approach also offers theoretical guarantees, including an asymptotic upper bound on post-calibration bias that decreases with data size, derived from OT theory.
Empirical results show that our calibration method effectively reduces DP score bias across various benchmarks and matching methods. However, it does not address EOD score bias, which depends on true labels. To tackle this, we introduce conditional calibration, a variant that adjusts scores based on both the predicted score and the estimated label. This method significantly reduces EOD score bias, as demonstrated through extensive evaluations on standard benchmarks.
@@
This paper is organized as follows. Section II presents background on record matching, notation, and existing fairness metrics. Section III introduces our proposed fairness metric and formalizes the problem. Sections IV and V describe two bias reduction algorithms. Section VI reports experimental results. Section VII reviews related and Section VIII concludes with a summary and future directions.
II Preliminaries
@@
II-A Record Matching
Given two relations, $R_{1}$ and $R_{2}$ with schemas $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$, record matching aims to identify a subset $E$ of record pairs in $R_{1}\times R_{2}$ that refer to the same real-world entities. These are called equivalent pairs, while all others are non-equivalent. Without loss of generality and for simplicity, we assume $R_{1}=R_{2}$.
A record matcher, or simply a matcher, is a binary classifier $f:\mathcal{R}^{2}\rightarrow\{0,1\}$ that assigns a label of $1$ (match) or 00 (non-match) to pairs of records from a relation $R$ with schema $\mathcal{S}$. The goal is to correctly label equivalent pairs as matches and non-equivalent pairs as non-matches.
A matcher $f$ is often defined using a matching score function $s:\mathcal{R}^{2}\to[0,1]$ and a threshold $\theta\in[0,1]$. It labels a pair $p$ as a match if $s(p)\geq\theta$, i.e., $f(p)=\mathbbm{1}_{s(p)\geq\theta}$.
[…]
@@ II-A Record Matching
$A$ Binary sensitive (protected) attribute
@@
$P(X)$ Probability distributions or random variable $X$
@@
TABLE I: Summary of notation and symbols.
II-B Fairness Measure in Record Matching
To study fairness in record matching, we assume records contain a sensitive attribute $A\in\mathcal{S}$, such as gender or ethnicity, which defines majority and minority groups. For simplicity, we consider $A$ to be binary, with $Dom(A)=\{a,b\}$, where records with value $a$ belong to the minority group (e.g., female) and those with $b$ to the majority group (e.g., male). A record pair $p=(r_{1},r_{2})$ is considered a minority pair, denoted by $p[A]=a$, if either record belongs to the minority group, i.e., if $r_{1}[A]=a$ or $r_{2}[A]=a$; otherwise, it is a majority pair ($p[A]=b$). This definition ensures that any pair involving a minority record is treated as a minority pair, reflecting that mismatches may disproportionately affect minority records.
II-B1 Fairness in Binary Matching
Fairness in record matching, treated as a binary classification task, has been extensively studied in [8]. Several fairness definitions are applied to assess bias, following principles similar to those used in general classification tasks. These definitions examine the relationship between the sensitive attribute $p[A]$ and the matcher’s output $f(p)$ for any record pair $p$. Each definition compares a performance metric $\Phi$, such as TPR, FPR, or PR, across sensitive groups, denoted $\Phi_{g}$ for $g\in\{a,b\}$. The corresponding fairness metric, or bias, is denoted as $\textit{bias}(f,\Phi)$.
Demographic parity, or statistical parity, requires the positive rate $\Phi=\text{PR}$ to be independent of the sensitive attribute, i.e., $\hat{Y}\mbox{$\perp\!\!\!\perp$}A$. This implies that the probability of matching is the same across groups, regardless of true labels. Formally, $\text{PR}_{a}=\text{PR}_{b}$, where $\text{PR}_{g}=P(\hat{Y}=1\mid A=g)$. The associated fairness metric is the demographic parity difference, given by $\textit{bias}(f,\text{PR})=|\text{PR}_{a}-\text{PR}_{b}|$.
Demographic parity ignores ground truth equivalence. In contrast, definitions such as equalized odds and equal opportunity condition on equivalence to evaluate fairness, requiring $\hat{Y}\mbox{$\perp\!\!\!\perp$}A\mid Y$. Equal opportunity (EO), also known as true positive rate parity, requires $\text{TPR}_{a}=\text{TPR}_{b}$, where $\text{TPR}_{g}=P(\hat{Y}=1\mid A=g,Y=1)$. The corresponding fairness metric is $\textit{bias}(f,\text{TPR})=|\text{TPR}_{a}-\text{TPR}_{b}|$.
Equalized odds (EOD) extends this further by requiring both TPR and FPR to be equal across groups. A matcher satisfies EOD if $\text{TPR}_{a}=\text{TPR}_{b}$ and $\text{FPR}_{a}=\text{FPR}_{b}$, where $\text{FPR}_{g}=P(\hat{Y}=1\mid A=g,Y=0)$. EOD captures fairness for both positive and negative classes. The fairness metric based on EOD is $\textit{bias}(f,\text{TPR}+\text{FPR})=|\text{TPR}_{a}-\text{TPR}_{b}|+|\text{FPR}_{a}-\text{FPR}_{b}|$.
II-C Wasserstein Distance and Barycenter
@@
III Problem Definition
As discussed in Section II-B, existing fairness definitions are either threshold-specific or, in the case of AUC-based, threshold-independent but potentially misleading. To address these limitations, we introduce a new fairness metric for matching scores, extending binary classification fairness measures (see Section II-B1) to evaluate biases in a score function $s$ across all thresholds.
@@ III Problem Definition
Definition 1 generalizes fairness notions such as DP, EO, and EOD to score functions. For example, setting $\Phi=\text{PR}$ recovers DP for $s$, referred to as score DP. Similarly, $\Phi=\text{TPR}$ corresponds to score EO, and using both $\Phi=\text{TPR}$ and $\Phi=\text{FPR}$ yields EOD. For any unfair score function, the value of $\textit{bias}(s,\Phi)$ quantifies the score bias. For example, $\textit{bias}(s,\text{PR})$ measures the DP score bias, $\textit{bias}(s,\text{TPR})$ captures the EO score bias, and $\textit{bias}(s,\text{TPR})+\textit{bias}(s,\text{FPR})$ represents the EOD score bias. When clear from context, we omit the term “score” and simply refer to these as DP, EO, or EOD biases.
Prior work [8, 5] has shown that many existing record matching methods produce biased matching scores. In Section VI-B1, we demonstrate through experiments on standard benchmarks that several state-of-the-art techniques exhibit substantial bias. To address this, we define the problem of generating fair matching scores and propose a post-processing solution to adjust scores from existing methods.
@@
IV Score Calibration Using Barycenter
@@ IV-A Algorithm Description and Analysis
The quantile-based method calculates the barycenter quantile function, ~$\tilde{\mathbb{Q}}(p)$, as a weighted average of the quantile functions of the original distributions:
~$\tilde{\mathbb{Q}}(p)=\alpha\mathbb{Q}_{a}(p)+(1-\alpha)\mathbb{Q}_{b}(p),$
@@
1 /* Initialize group score lists */
2 $\text{scores}_{b}\leftarrow[]$; $\text{scores}_{a}\leftarrow[]$;

### v3

@@ Abstract
Record matching models typically output a real-valued matching score that is later consumed through thresholding, ranking, or human review. While fairness in record matching has mostly been assessed using binary decisions at a fixed threshold, such evaluations can miss systematic disparities in the entire score distribution and can yield conclusions that change with the chosen threshold.
We introduce a threshold-independent notion of score bias that extends standard group-fairness criteria—demographic parity (DP), equal opportunity (EO), and equalized odds (EOD)—from binary outputs to score functions by integrating group-wise metric gaps over all thresholds. Using this metric, we empirically show that several state-of-the-art deep matchers can exhibit substantial score bias even when appearing fair at commonly used thresholds.
To mitigate these disparities without retraining the underlying matcher, we propose two model-agnostic post-processing methods that only require score evaluations on an (unlabeled) calibration set. Calib targets DP by aligning minority/majority score distributions to a common Wasserstein barycenter via a quantile-based optimal-transport map, with finite-sample guarantees on both residual DP bias and score distortion. C-Calib extends this idea to label-dependent notions (EO/EOD) by performing barycenter alignment conditionally on an estimated label, and we characterize how its guarantees depend on both sample size and label-estimation error.
Experiments on standard record-matching benchmarks and multiple neural matchers confirm that Calib and C-Calib substantially reduce score bias with minimal loss in accuracy.
keywords:
Record Matching , Fairness , Wasserstein Barycenter , Calibration , Threshold-Independent
journal: Data & Knowledge Engineering
1 Introduction
@@
From a technical perspective, matching scores support the development of more flexible and advanced methods. For example, active learning approaches can use scores to identify uncertain cases—those with values near the decision threshold—and prioritize them for manual labeling to improve the model [20, 25]. This idea also extends to human-in-the-loop systems, where scores help determine which record pairs should be reviewed by a human. Instead of treating all pairs equally, the system can focus on those that are harder to classify automatically [33, 15]. Another technical reason for using scores instead of a binary outcome is that many state-of-the-art matching techniques, especially deep learning models, naturally produce scores through their architecture—for example, by applying a sigmoid activation to the output of a final linear layer, resulting in a value between 0 and 1 that reflects the predicted match likelihood. Because of these advantages, matching scores are more widely used than hard binary decisions in both research and practical systems.
@@
Example 1.
Figure 1b compares the ROC curves of the matcher for the minority and majority groups. Each ROC curve summarizes performance across all thresholds by plotting TPR versus FPR, and the area under the curve (AUC) is a common threshold-independent measure of score quality. A standard fairness check is to compare AUC across groups.
In this example, however, the group AUC values are almost identical even though the ROC curves differ: one group has higher TPR in some regions of FPR and lower TPR in others. These differences can matter at the thresholds used in practice, but they can cancel out when aggregated into a single AUC number. This illustrates why AUC-based, threshold-independent comparisons can still miss meaningful score bias. $\blacksquare$
@@
Fairness in matching scores is closely related to fairness in both regression and ranking, since all three settings use a real-valued score that may exhibit systematic differences across protected groups.
In fair regression, the output is a continuous prediction, and many approaches aim to reduce group disparities by constraining or post-processing these numerical outputs [16, 1]. A particularly relevant line of work treats predictions as samples from group-wise distributions and then aligns these distributions via post-processing (e.g., using optimal transport and Wasserstein barycenters) [10]. This is closely aligned with our approach: we view matching scores for the minority and majority as two empirical distributions and post-process them to a common target distribution (a barycenter), thereby reducing threshold-independent score disparities while keeping score distortion small.
Fair ranking is also relevant because ranking pipelines rely on scores that are then consumed through ordering (e.g., top-$k$ selection). Many fair-ranking methods seek to enforce group-level representation or exposure constraints in the top-ranked results (e.g., [42, 7, 6]). These methods highlight how biases in scores can translate into biased downstream decisions, and they provide a complementary perspective to our focus on fairness of the scores themselves.
More broadly, there are many other techniques in both fair regression and fair ranking, including in-processing constraints and alternative post-processing strategies [16, 1, 42, 7, 6]. In this work, we primarily build on the distributional-alignment perspective from fair regression, since it naturally matches our goal of reducing disparities in the entire score distribution.
At the same time, record matching introduces important differences with fair regression and fair ranking. Matching scores are typically bounded in $[0,1]$, often concentrated near 0 and 1, and are used through thresholding, ranking, and human-in-the-loop review. More importantly, record matching is a setting where both the score and the induced binary decision are meaningful outcomes, so fairness notions naturally include label-dependent criteria such as equal opportunity (EO) and equalized odds (EOD). These criteria require aligning error rates (e.g., TPR/FPR) across groups and therefore depend on the (unknown) true match/non-match label. This label dependence is a main technical difference from most ranking and regression settings (and from many of their fairness objectives), where the fairness constraint is typically defined without conditioning on ground-truth labels. Thus, while distributional alignment ideas can directly target threshold-independent score disparity, extending them to EO/EOD is nontrivial when labels are unavailable at calibration time. This motivates our conditional calibration approach, which conditions distributional alignment on an estimated label so it can target EO/EOD in settings where true labels are unavailable.
Our contributions are as follows.
@@
(1) Score-bias formulation for record matching. We introduce a threshold-independent bias measure that extends standard notions such as demographic parity (DP) and equalized odds (EOD) from binary outcomes to score functions. It captures cumulative performance differences between groups over the full threshold range. For example, in Figure 1a, the DP score bias corresponds to the area between group-wise TPR curves across thresholds.
(2) Barycenter-based score calibration for DP. Building on distributional calibration ideas from the fair regression literature, we propose a model-agnostic post-processing algorithm that maps group-wise score distributions to a common Wasserstein barycenter [10, 27]. This reduces DP score bias while keeping score distortion small; moreover, our theoretical guarantees for the DP case build directly on theory developed in the fair regression literature.
(3) Conditional calibration for label-dependent notions (EO/EOD). We introduce conditional calibration, which applies barycenter-based alignment within strata defined by an estimated label (since ground-truth labels may be unavailable at calibration time). This extends score calibration beyond DP to address label-dependent disparities such as EO and EOD, and we extend the theoretical analysis to this setting by characterizing how EO/EOD fairness depends on both sample size and label-estimation error.
(4) Empirical evaluation. Experiments across standard record-matching benchmarks and state-of-the-art matchers show that the proposed methods substantially reduce score bias with minimal impact on accuracy.
@@
This paper is organized as follows. Section 2 presents background on record matching, notation, and existing fairness metrics. Section 3 introduces our proposed fairness metric and formalizes the problem. Sections 4 and 5 describe two bias reduction algorithms. Section 6 reports experimental results. Section 7 reviews related work and Section 8 concludes with a summary and future directions.
2 Preliminaries
@@
2.1 Record Matching
Given two relations, $R_{1}$ and $R_{2}$ with schemas $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$, record matching aims to identify a subset $E$ of record pairs in $R_{1}\times R_{2}$ that refer to the same real-world entities. These are called equivalent pairs, while all others are non-equivalent. Without loss of generality and for simplicity, we assume $R_{1}=R_{2}$. A record matcher, or simply a matcher, is a binary classifier $f\mathrel{\mathop{\ordinarycolon}}\mathcal{R}^{2}\rightarrow\{0,1\}$ that assigns a label of $1$ (match) or 0 (non-match) to pairs of records from a relation $R$ with schema $\mathcal{S}$. The goal is to correctly label equivalent pairs as matches and non-equivalent pairs as non-matches. A matcher $f$ is often defined using a matching score function $s\mathrel{\mathop{\ordinarycolon}}\mathcal{R}^{2}\to[0,1]$ and a threshold $\theta\in[0,1]$. It labels a pair $p$ as a match if $s(p)\geq\theta$, i.e., $f(p)=\mathbbm{1}_{s(p)\geq\theta}$. Performance metrics for $f$ are formally defined with respect to a joint probability distribution $P$ over a random variable $X$ (representing pairs of records from $\mathcal{S}$) and a binary variable $Y$ indicating the true label. The positive rate (PR), true positive rate (TPR), and false positive rate (FPR) for $s$ and $\theta$ are defined as $\text{PR}(s,\theta)=P(s(X)\geq\theta)$, $\text{TPR}(s,\theta)=P(s(X)\geq\theta\mid Y=1)$, and $\text{FPR}(s,\theta)=P(s(X)\geq\theta\mid Y=0)$. We use $\hat{Y}$ to refer to $f(X)$ and, when clear from context, write $A_{i}$ instead of $A$ and $X[A_{i}]$ for attribute values. Other metrics such as precision, recall, and F1 score are defined similarly. The AUC for $s$ can be written as $\int_{0}^{1}\text{TPR}(s,\text{FPR}^{-1}(s,u))\,du$.
@@ 2.1 Record Matching
AA Binary sensitive (protected) attribute
@@
$P(X)$ Probability distributions with random variable $X$
@@
Table 1: Summary of notation and symbols.
2.2 Fairness Measure in Record Matching
To study fairness in record matching, we assume records contain a sensitive attribute $A\in\mathcal{S}$, such as gender or ethnicity, which defines majority and minority groups. For simplicity, we consider $A$ to be binary, with $Dom(A)=\{a,b\}$, where records with value $a$ belong to the minority group (e.g., female) and those with $b$ to the majority group (e.g., male). A record pair $p=(r_{1},r_{2})$ is considered a minority pair, denoted by $p[A]=a$, if either record belongs to the minority group, i.e., if $r_{1}[A]=a$ or $r_{2}[A]=a$; otherwise, $r_{1}[A]=b$ and $r_{2}[A]=b$ and it is a majority pair, which we denote by $p[A]=b$. This definition ensures that any pair involving a minority record is treated as a minority pair, reflecting that mismatches may disproportionately affect minority records.
2.2.1 Fairness in Binary Matching
Fairness in record matching, treated as a binary classification task, has been extensively studied in [34]. Several fairness definitions are applied to assess bias, following principles similar to those used in general classification tasks. These definitions examine the relationship between the sensitive attribute $p[A]$ and the matcher’s output $f(p)$ for any record pair $p$. When a matcher is defined by a score function $s$ and a threshold $\theta$, these fairness metrics are threshold-dependent; we make this explicit using a subscript $\theta$ (e.g., $\text{DP}_{\theta}$).
Each definition compares a performance metric $\Phi$, such as TPR, FPR, or PR, across sensitive groups, denoted $\Phi_{g}$ for $g\in\{a,b\}$. Demographic parity, or statistical parity, requires the positive rate (PR) to be independent of the sensitive attribute, i.e., $\hat{Y}\mbox{$\perp\!\!\!\perp$}A$. Formally, $\text{PR}_{a}(s,\theta)=\text{PR}_{b}(s,\theta)$, where $\text{PR}_{g}(s,\theta)=P(\hat{Y}=1\mid A=g)$ and $\hat{Y}=\mathbbm{1}_{s(X)\geq\theta}$. The corresponding demographic parity difference is $\text{DP}_{\theta}=|\text{PR}_{a}(s,\theta)-\text{PR}_{b}(s,\theta)|$.
In contrast, definitions such as equalized odds and equal opportunity condition on equivalence to evaluate fairness, requiring $\hat{Y}\mbox{$\perp\!\!\!\perp$}A\mid Y$. Equal opportunity (EO) requires $\text{TPR}_{a}(s,\theta)=\text{TPR}_{b}(s,\theta)$, where $\text{TPR}_{g}(s,\theta)=P(\hat{Y}=1\mid A=g,Y=1)$. The corresponding difference is $\text{EO}_{\theta}=|\text{TPR}_{a}(s,\theta)-\text{TPR}_{b}(s,\theta)|$.
Equalized odds (EOD) extends this further by requiring both TPR and FPR to be equal across groups. A matcher satisfies EOD at threshold $\theta$ if $\text{TPR}_{a}(s,\theta)=\text{TPR}_{b}(s,\theta)$ and $\text{FPR}_{a}(s,\theta)=\text{FPR}_{b}(s,\theta)$, where $\text{FPR}_{g}(s,\theta)=P(\hat{Y}=1\mid A=g,Y=0)$. We quantify this via $\text{EOD}_{\theta}=|\text{TPR}_{a}(s,\theta)-\text{TPR}_{b}(s,\theta)|+|\text{FPR}_{a}(s,\theta)-\text{FPR}_{b}(s,\theta)|$.
2.3 Wasserstein Distance and Barycenter
@@
3 Problem Definition
As discussed in Section 2.2, existing fairness definitions are either threshold-specific or, in the case of AUC-based, threshold-independent but potentially misleading. To address these limitations, we introduce a new fairness metric for matching scores, extending binary classification fairness measures (see Section 2.2.1) to evaluate biases in a score function $s$ across all thresholds.
@@ 3 Problem Definition
Definition 1 generalizes fairness notions such as DP, EO, and EOD to score functions. For example, setting $\Phi=\text{PR}$ recovers DP for $s$, referred to as score DP. Similarly, $\Phi=\text{TPR}$ corresponds to score EO, and using both $\Phi=\text{TPR}$ and $\Phi=\text{FPR}$ yields EOD. For any unfair score function, the value of $\textit{bias}(s,\Phi)$ quantifies the score bias. For example, $\textit{bias}(s,\text{PR})$ measures the DP score bias, $\textit{bias}(s,\text{TPR})$ captures the EO score bias, and $\textit{bias}(s,\text{TPR})+\textit{bias}(s,\text{FPR})$ represents the EOD score bias. When clear from context, we omit the term “score” and simply refer to these as DP, EO, or EOD biases; we reserve subscripts (e.g., $\text{DP}_{\theta}$, $\text{EO}_{\theta}$, $\text{EOD}_{\theta}$) for threshold-specific binary fairness metrics.
Prior work [34, 36] has shown that many existing record matching methods produce biased matching scores. In Section 6.2.1, we demonstrate through experiments on standard benchmarks that several state-of-the-art techniques exhibit substantial bias. To address this, we define the problem of generating fair matching scores and propose a post-processing solution to adjust scores from existing methods.
@@
4 Score Calibration Using Barycenter
@@ 4.1 Calib Algorithm
The quantile-based method calculates the barycenter quantile function, $\tilde{\mathbb{Q}}(p)$, as a weighted average of the quantile functions of the original distributions:
$\tilde{\mathbb{Q}}(p)=\alpha\mathbb{Q}_{a}(p)+(1-\alpha)\mathbb{Q}_{b}(p),$
@@
1
[…]